OpenStudy (anonymous):

what is the factor of x^3-8y^3?

OpenStudy (anonymous):

(x-y)(x^2+xy+y^2)

mathslover (mathslover):

\[\Huge{\mathbb{Welcome}\textbf{To }\mathbb{Open}\textbf{Study}}\] \[\huge{x^3-y^3=(x-y)(x^2+xy+y^2)}\]]

OpenStudy (ash2326):

@roanne did you understand?

mathslover (mathslover):

please reply if u got it @roanne